# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v4

@@ Abstract
The significant features identified in a representative subset of the dataset during the learning process of an artificial intelligence model are referred to as a ’global’ explanation. Three-dimensional (3D) global explanations are crucial in neuroimaging where a complex representational space demands more than basic two-dimensional interpretations. Curently, studies in the literature lack accurate, low-complexity, and 3D global explanations in neuroimaging and beyond. To fill this gap, we develop a novel explainable artificial intelligence (XAI) 3D-Framework that provides robust, faithful, and low-complexity global explanations. We evaluated our framework on various 3D deep learning networks trained, validated, and tested on a well-annotated cohort of 596 MRI images. The focus of detection was on the presence or absence of the paracingulate sulcus, a highly variable feature of brain topology associated with symptoms of psychosis. Our proposed 3D-Framework outperformed traditional XAI methods in terms of faithfulness for global explanations. As a result, these explanations uncovered new patterns that not only enhance the credibility and reliability of the training process but also reveal the broader developmental landscape of the human cortex. Our XAI 3D-Framework proposes for the first time, a way to utilize global explanations to discover the context in which detection of specific features are embedded, opening our understanding of normative brain development and atypical trajectories that can lead to the emergence of mental illness.
@@ 1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. Recently, the study by Mamalakis et al. (2024) introduced the necessity of explanations in artificial intelligence (AI) healthcare applications, categorizing them into four types: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions, the stability of the evaluation protocol, and the dimensionality of the problem.
In neuroscience AI applications, there commonly exists significant variability in evaluation protocols and decision-making among experts. The inherent high uncertainty creates a greater need for explainability (as the framework falls in the non-explainable category; Mamalakis et al. (2024)). To this end, applications usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, and this is where AI has the potential to offer insights and stabilize the validity of key aspects of the evaluation protocols (Mitchell et al. (2023)). This is particularly true for classification tasks where the key features of a disease are not yet firmly established (matching the new-patterns discovery category; Mamalakis et al. (2024)).
@@
While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some sulci present a more complex computational hurdle (Borne et al. (2020)). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental period with information that may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues (Jiang et al. (2021)). In particular, the paracingulate sulcus (PCS), a highly variable sulcus, is associated with cognitive performance and hallucinations in schizophrenia (Fornito et al. (2006), Garrison et al. (2015), Gay et al. (2021), Ćurčić-Blake et al. (2023)). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in brain magnetic resonance imaging (MRI) and creating frameworks that can provide both local and global explanations would allow for a more systematic characterization and a more comprehensive understanding of whole-brain correlates related to its presence. This, in turn, would allow for both more precise and holistic assessments of its functional relevance and impact on brain functions (Simons et al. (2017)).
@@
However, an important challenge arises, particularly in neuroimaging, where the high representation dimensionality in AI applications, the high anatomical complexity, the involvement of unstable evaluation protocols, and the intricate yet uncertain expert annotations make these tasks challenging to explain and evaluate. Ordinary 2D local or global XAI methods fall short and may even be incorrect in some cases (Mamalakis et al. (2024)). Thus, there is a need for 3D explanation frameworks that can surpass traditional XAI methods and deliver more reliable global explanations.
Contribution. For the first time, we propose a comprehensive framework to validate and explain deep learning models in a way that reveals novel insights into pattern learning, thereby enhancing the credibility and reliability of these networks. To achieve this, we developed a novel XAI 3D-Framework designed to assess the ability of deep learning networks to learn patterns in a binary classification task to detect the presence or absence of the PCS, a highly variable sulcus related to reality monitoring and psychotic conditions.
We applied two distinct local XAI methods, GradCam and SHAP, in 3D space to the entire dataset used in the binary classification task. A dimensionality reduction technique was then employed to extract global explanations from these XAI methods, as well as statistical patterns from the 3D brain inputs. By combining these two XAI methods with the statistical patterns in various configurations we outperformed traditional XAI methods in global explainability while revealing the learning patterns within the network.
For the detection of the PCS presence we developed two different 3D deep learning networks: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with variations in backbone options (2CNN-3D-MHL and simple-3D-MHL). The networks utilized 3D brain inputs derived from preprocessing structural MRI scans, which included grey-white surface boundaries and sulcal skeletons from both the right and left hemispheres. We used a well-annotated cohort of 596 subjects from the TOP-OSLO study cohort (Mørch-Johnsen et al. (2017)) to train, validate, and test our deep learning networks, with a split of 70% for training, 20% for validation, and 10% for testing.
The proposed 3D framework surpassed traditional XAI methods in terms of faithfulness for global explanation. Importantly, our framework effectively identified significant sub-regions by combining explainable features with statistical characteristics from the evaluation dataset. This approach offers a robust means to determine the patterns that contribute to a network’s decisions, thereby providing greater transparency, trust and further exploration of the outcomes of deep learning models used in neuroscience applications.
2 Related work
2.1 The application: Sulcal pattern studies
Cortical folding, which develops in the perinatal period (i.e. in the last few months of gestation and the first few months after birth) results in an inter-individual variability that is often eschewed in populational studies. Interest in the variability of sulcal patterns can be motivated by different reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of mechanisms at play during neurodevelopment, and a search for anatomo-functional correlates. Links to brain function can also be addressed to investigate healthy functional variability (e.g. cingulate folding pattern and functional connectivity Fedeli et al. (2022)), or relationships between cortical folding pattern and pathological outcomes (e.g. paracingulate folding and hallucinations in schizophrenia Rollins et al. (2020)). The study of cortical folding variability can be addressed through global methods, with whole brain or regional considerations of generic sulcal parameters (e.g. gyrification index or sulcal pits), but most refined investigations of sulcal patterns require the capability for a focus on specific sulci, hence the need for automated sulcal recognition methods. A number of different methods have been developed for automated sulcal labelling in a general setting (for review, see Mangin et al. (2015)). Yet, to the best of our knowledge, no method is currently implemented to automatically label the PCS in a 3D approach (a 2D approach for a specific MRI slice has been proposed Yang et al. (2019)). The fact that the PCS is generally omitted in whole-brain labelling is related to the way in which it is anatomically specified; not only by its location in the brain, but also by its orientation (parallel to the cingulate sulcus). Hence, even newer sulcal labelling implementations do not identify the PCS within the folding of the medial frontal cortex (Borne et al. (2020)). Nevertheless, investigating the whole-brain anatomical correlates of sulcal variability incurred by the presence of the PCS is potentially important to understand some of the most common and severe symptoms of psychosis. Automatic detection of the PCS is in itself useful for large dataset exploration, especially since it has previously been related to functional variability in reality monitoring (Simons et al. (2017)). Additionally and unprecedentedly, adding an explainability component allows us to interrogate the anatomical covariates of the PCS gaining a broader vision of the functional mechanisms in which it may have a role.
2.2 Explainable methods
@@ 3 Methods and Background
In this section, we describe the different 3D classifiers used to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we develop a novel XAI 3D-Framework for the non-explainable and new-patterns discovery tasks (Mamalakis et al. (2024)). We utilize two distinct explanation methods, belonging to the post-hoc family SHAP and GradCam, expanded in three-dimensional space. We concatenate the outcomes obtained from the two explainable methods with the results derived from a statistical feature extraction model. The statistical model is a transparent dimensionality reduction algorithm applied to the input data of the validation cohort. This whole approach is intended to mitigate any potential biases and enhance the robustness of the pattern learning discovery.
@@ 3.3 The XAI 3D-Framework
The main purpose of the XAI 3D-Framework is to provide global explanations that can lead to the discovery of new patterns. To achieve this, we used two XAI methods, GradCam and SHAP in 3D space to extract local explanations from the deep learning classifiers (simple-3D-CNN, simple-3D-MHL; Fig. 2 b.). Concurrently, we assigned faithfulness and complexity scores to evaluate the quality of the local explanations ensuring that their effectiveness reached an acceptable threshold (see Supplementary material Table 2). Following that, we used Principal Component Analysis (PCA; (Pearson (1901)) to derive global explanations from the two XAI methods. PCA dimensionality reduction was applied to both the sulcal skeleton and the grey-white surface 3D brain inputs (analytical description of the 3D inputs please advice sub-section 4.2) to identify a variety of global feature importance (feature-shape; see Fig. 3). We experimented with a range of different numbers of PCA components to find the optimal configuration in terms of time efficiency and the proportion of cohort variance captured ($>80.0\%$). The optimal solution was found with six components.
Based on preliminary research, we found that the global explanations derived from each of the two XAI methods (GradCam and SHAP) yielded better results in terms of faithfulness and complexity when we averaged the six components of the PCA analysis using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs total-SHAP, and PCA0-GradCam vs total-GradCam). Consequently, to generate the final global explanation represented from the GradCam (total-GradCam) and SHAP (total-SHAP), we applied the weighted average to the six PCA components (see Fig. 3). The same approach was applied in the feature-shape domain to determine the global statistical feature importance of the sulcal skeleton and the grey-white surfaces in the 3D brain inputs (total-Shape; see Fig. 3). The weighted average formulation is defined as follows:
@@
3.4 The global explanation of 3D-Framework
To compute the global explanation for the 3D framework, we first manually aligned and rescaled the two global XAI explanations (total-SHAP, total-GradCam) to the global statistical features (total-Shape). Then, we used equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. We conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which are fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey-white surface inputs of the total dataset, considering both the right and left hemispheres.
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. Identifying the significant features of the networks explanations provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey-white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern associated with determining the presence or absence of the PCS. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas.
4 Experiments

### v5

@@ Abstract
The significant features identified in a representative subset of the dataset during the learning process of an artificial intelligence model are referred to as a ’global’ explanation. Three-dimensional (3D) global explanations are crucial in neuroimaging, where a complex representational space demands more than basic two-dimensional interpretations. However, current studies in the literature often lack the accuracy, comprehensibility, and 3D global explanations needed in neuroimaging and beyond. To address this gap, we developed an explainable artificial intelligence (XAI) 3D-Framework capable of providing accurate, low-complexity global explanations. We evaluated the framework using various 3D deep learning models trained on a well-annotated cohort of 596 structural MRIs. The binary classification task focused on detecting the presence or absence of the paracingulate sulcus (PCS), a highly variable brain structure associated with psychosis. Our framework integrates statistical features (Shape) and XAI methods (GradCam and SHAP) with dimensionality reduction, ensuring that explanations reflect both model learning and cohort-specific variability. By combining Shape, GradCam, and SHAP, our framework reduces inter-method variability, enhancing the faithfulness and reliability of global explanations. These robust explanations facilitated the identification of critical sub-regions, including the posterior temporal and internal parietal regions, as well as the cingulate region and thalamus, suggesting potential genetic or developmental influences. For the first time, this XAI 3D-Framework leverages global explanations to uncover the broader developmental context of specific cortical features. This approach advances the fields of deep learning and neuroscience by offering insights into normative brain development and atypical trajectories linked to mental illness, paving the way for more reliable and interpretable AI applications in neuroimaging.
@@ 1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. Recently, the study by Mamalakis et al. (2024a) introduced the necessity of explanations in artificial intelligence (AI) healthcare applications, categorizing them into four types: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions, the stability of the evaluation protocol, and the dimensionality of the problem.
In neuroscience AI applications, there commonly exists significant variability in evaluation protocols and decision-making among experts. The inherent high uncertainty creates a greater need for explainability (as the framework falls in the non-explainable category; Mamalakis et al. (2024a)). To this end, applications usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, and this is where AI has the potential to offer insights and stabilize the validity of key aspects of the evaluation protocols (Mitchell et al. (2023)). This is particularly true for classification tasks where the key features of a disease are not yet firmly established (matching the new-patterns discovery category; Mamalakis et al. (2024a)).
@@
While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some sulci present a more complex computational hurdle (Borne et al. (2020)). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental period with information that may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues (Jiang et al. (2021)). In particular, the morphology of the paracingulate sulcus (PCS), a highly variable sulcus, is associated with cognitive performance and hallucinations in schizophrenia (Fornito et al. (2006), Garrison et al. (2015), Gay et al. (2021), Ćurčić-Blake et al. (2023)). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in brain magnetic resonance imaging (MRI) and creating frameworks that can provide both local and global explanations would allow for a more systematic characterization and a more comprehensive understanding of whole-brain correlates related to its presence. This, in turn, would allow for both more precise and holistic assessments of its functional relevance and impact on brain functions (Simons et al. (2017)).
@@
However, a significant challenge arises in neuroimaging due to the high representational dimensionality of AI applications, the anatomical complexity of the brain, the use of unstable evaluation protocols, and the intricate yet uncertain nature of expert annotations. These factors make tasks in this domain particularly difficult to explain and evaluate. Conventional 2D local or global XAI methods often fall short and, in some cases, may even produce misleading explanations (Mamalakis et al. (2024a)). Additionally, as highlighted in (Mamalakis et al., 2024b), the issue of inter-method variability—where different XAI methods emphasize different features as important—can significantly erode trust in AI within the scientific community. These challenges underscore the urgent need for 3D explanation frameworks that can surpass the limitations of traditional XAI methods and provide more reliable global explanations.
1.1 Aim and contribution of the study
This study introduces, for the first time, a comprehensive framework to validate and explain deep learning models, providing transformative insights into pattern learning while establishing a new standard of credibility and reliability for such networks. The proposed XAI 3D-Framework tackles the intricate challenges of explainability in neuroimaging, enabling the detection and interpretation of complex patterns related to the presence or absence of the paracingulate sulcus (PCS), a highly variable cortical structure associated with reality monitoring and psychotic conditions.
To achieve this, we developed an innovative methodology that employs two complementary local XAI methods, GradCam and SHAP, extended into 3D space to analyze the entire dataset used in the binary classification task. These methods are integrated with statistical patterns derived from the 3D brain inputs (Shape) via dimensionality reduction, producing global explanations that outperform traditional XAI methods in both interpretability and faithfulness. By combining these complementary approaches, our framework not only reveals the underlying learning patterns within the network but also significantly enhances the accuracy and clarity of the results. For the PCS classification task, we implemented two distinct 3D deep learning architectures: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with diverse backbone options (2CNN-3D-MHL and simple-3D-MHL). These networks utilized high-resolution 3D brain inputs, including grey-white surface boundaries and sulcal skeletons from both hemispheres. Leveraging a well-annotated cohort of 596 subjects from the TOP-OSLO study (Mørch-Johnsen et al. (2017)), we trained, validated, and tested these networks, employing a 70%-20%-10% data split.
Our framework introduces several pivotal innovations to the domain of explainable AI in neuroimaging: (i) By integrating statistical features (Shape) that are correlated with reduced dimensionality information, the framework ensures that the discovered patterns are not only grounded in the AI model’s learning but also reflect cohort-specific variability. This dual-layered approach bridges statistical data and model-derived insights, enabling a deeper and more contextually relevant understanding of the results. (ii) The extension of established XAI methods, such as GradCam and SHAP, into the 3D domain addresses the critical need for 3D explanations in neuroimaging applications. Conventional 2D local or global XAI methods often fall short and, in some cases, may even produce misleading explanations. (iii) The use of GradCam and SHAP in tandem reduces inter-method variability and bolsters the reliability of the explanations, setting a new benchmark for trustworthy AI applications. The proposed multi-method framework delivers robust and actionable insights, particularly in complex neuroimaging tasks such as cortical morphology studies.
Notably, the XAI 3D-Framework demonstrated superior performance compared to traditional XAI methods in terms of faithfulness for global explanations, successfully identifying significant sub-regions of an atlas brain (the ICBM 2009a Nonlinear Asymmetric atlas, Fonov et al. (2011, 2009)). This capability provides a transparent and reliable mechanism to trace the patterns driving network decisions, enhancing trust and enabling deeper exploration of deep learning model outputs in neuroscience. By combining methodological rigor with practical innovation, this framework opens new avenues for understanding and interpreting brain structure-function relationships, making it a foundational tool for advancing both research and clinical applications in neuroimaging.
1.2 Related work
1.2.1 The application: Sulcal pattern studies
Cortical folding, which develops during the perinatal period (i.e., in the last few months of gestation and the first few months after birth), results in significant inter-individual variability often overlooked in population studies. Understanding the variability of sulcal patterns is critical for multiple reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of neurodevelopmental mechanisms, and the search for anatomo-functional correlates. These patterns hold potential for investigating healthy functional variability (e.g., the relationship between cingulate folding patterns and functional connectivity Fedeli et al. (2022)) and pathological outcomes (e.g., paracingulate folding linked to hallucinations in schizophrenia Rollins et al. (2020)).
While the study of cortical folding variability has been approached through global methods—considering whole-brain or regional sulcal parameters such as gyrification index or sulcal pits—finer investigations often require a focus on specific sulci. This underscores the need for automated sulcal recognition methods. Although several techniques have been developed for general sulcal labeling (reviewed in Mangin et al. (2015)), to the best of our knowledge, no current method can automatically label the PCS in a 3D approach; most rely on 2D analyses of specific MRI slices (Yang et al. (2019)). The omission of PCS in whole-brain labeling stems from its complex anatomical specification, defined not only by its location but also by its orientation (parallel to the cingulate sulcus). Consequently, even newer labeling frameworks fail to identify the PCS within the medial frontal cortex (Borne et al. (2020)).
This gap is particularly critical given the potential importance of investigating whole-brain anatomical correlates of PCS variability in understanding severe symptoms of psychosis. Automatic detection of the PCS is valuable for large dataset exploration, especially given its links to functional variability in reality monitoring (Simons et al. (2017)). Moreover, incorporating an explainability component through advanced XAI frameworks is unprecedented and transformative. It allows not only for the identification of the PCS but also for uncovering new patterns in its anatomical covariates, offering insights into the functional mechanisms underlying its role. The integration of XAI with AI classification thus provides a robust platform for both enhancing interpretability and advancing our understanding of the broader neurodevelopmental and pathological contexts associated with the PCS.
1.2.2 Explainable methods
@@ 2 Methods and Background
In this study, we utilized different 3D classifiers to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we developed a novel XAI 3D framework for non-explainable and new-pattern discovery tasks (Mamalakis et al. (2024a)). We employed two distinct explanation methods from the post-hoc family, SHAP and GradCam, which were expanded into three-dimensional space. The outcomes from these two explainable methods (SHAP and GradCam) were concatenated with results derived from a statistical feature extraction model (Shape). The statistical model is a transparent dimensionality reduction algorithm applied to the input data of the validation cohort. This comprehensive approach aims to mitigate potential biases and enhance the robustness of pattern learning discovery.
@@ 2.3 The XAI 3D-Framework
The XAI 3D-Framework introduces a groundbreaking approach to generating global explanations that facilitate the discovery of new patterns in neuroimaging studies. Our proposed framework uniquely integrates statistical features derived from cohort data (Shape) with insights from two complementary explainability methods, GradCam and SHAP, in a three-dimensional space. By combining statistical information with model-driven learning, the framework provides a dual-layer understanding: one rooted in the cohort’s inherent variability and the other reflecting the AI model’s decision-making processes based on the classification task. Such integration not only enhances robustness but also minimizes inter-method variability and potential biases (Mamalakis et al. (2024b)), ensuring reliable and interpretable outcomes. The use of multiple methods, as opposed to relying on a single explainability approach, represents a significant step forward in delivering more comprehensive and trustworthy explanations, especially in challenging contexts like cortical morphology.
To achieve its goals, the framework employs GradCam and SHAP, two widely used explainability methods, expanded into three-dimensional space to extract local explanations from deep learning classifiers (simple-3D-CNN and simple-3D-MHL; Fig. 2 b.). Faithfulness and complexity scores are assigned to evaluate the quality of these local explanations, ensuring they meet predefined thresholds (see Supplementary Material Table 2). To generate global explanations, we apply Principal Component Analysis (PCA; (Pearson (1901))) to both the sulcal skeleton and grey-white surface 3D brain inputs, capturing a variety of global feature importance patterns (PCA-Shape; see Fig. 3). After testing different configurations, the optimal solution—capturing over 80% of the cohort variance—was achieved with six PCA components.
The global explanations derived from GradCam and SHAP yielded superior faithfulness and complexity results when averaged across the six PCA components using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam). These weighted averages formed the final global explanations (total-GradCam and total-SHAP). In parallel, the same approach was applied in the Shape domain to determine the global statistical feature importance of sulcal skeleton and grey-white surface inputs (total-Shape; see Fig. 3).
By integrating statistical and model-driven insights, the framework provides a robust platform for uncovering meaningful patterns while maintaining the reliability and interpretability essential for applications in neuroimaging. The weighted average formulation is defined as follows:
@@
2.4 The global explanation of the 3D-Framework
To compute the global explanation of the 3D framework, we first manually aligned and rescaled the two global XAI explanations (total-SHAP, total-GradCam) to the average of the six components of the PCA-Shape (total-Shape). Then, we used equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. The weight combination [0.85, 0.5, 0.1] was chosen after empirical testing of multiple combinations, to optimize faithfulness while maintaining reasonable complexity scores. Preliminary results showed that higher weights for total-Shape (e.g., 0.85) yielded superior faithfulness scores, reflecting the importance of feature importance in global explanations. This selection balances the trade-offs between faithfulness, complexity, and redundancy across the metrics. After we define the best combination, we conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which are fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey-white surface inputs of the total dataset, considering both the right and left hemispheres. In our specific study, the best combinations identified were 851 and 815 (see Supplementary Material Table 2).
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. Identifying the significant features of the networks explanations provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey-white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern associated with determining the presence or absence of the PCS. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas (the ICBM 2009a Nonlinear Asymmetric atlas, Fonov et al. (2011, 2009)).